# Towards the Training of Deeper Predictive Coding Neural Networks

### v2

@@ Appendix C Ablation Study
Figure 6: Layer-wise Energy Distribution and Accuracy of Forward Update in VGG7 and VGG10 on the CIFAR10 dataset. The colored lines represent the total energy of the individual layers of the model. The vertical lines represent the train and test accuracies of the model.
@@
In this section, we conducted ablation studies to evaluate the effectiveness of each proposed component. These experiments allowed us to quantify the individual contribution of each component and their synergistic effects when combined. The results are presented in Table 1.
Firstly, we observe that removing forward updates (using the standard $PC$ model with decaying/spiking Precision) causes significant accuracy degradation, which becomes more pronounced with increasing network depth. Simultaneously, the effectiveness of center nudging reappears when forward updates are removed, with its impact also strengthening in deeper networks as shown in Figure 4. This phenomenon confirms our earlier hypothesis: synaptic weight updates based on $x_{t}^{l}$ may introduce errors that accumulate across layers, leading to performance degradation in deep architectures.
Secondly, we observed that removing the decaying/spiking precision module leads to performance degradation in most cases. This effect is evident in deep models such as VGG7/10, where the energy distribution across layers becomes significantly imbalanced without the precision term. As shown in Figure 6 and Figure 5, in the VGG7 architecture with spiking precision and forward update, the first layer’s energy proportion is approximately $10^{-6}$, whereas in the model without spiking precision, this proportion drops dramatically to $10^{-12}$. Visualization of layer-wise energy distributions between decaying/spiking precision + forward update versus only forward update confirms that our proposed precision methods successfully balance energy distribution across layers, which contributes to improved model performance.
@@
Thirdly, we found that BatchNorm Freezing ($BF$) significantly enhances model performance when combined with our precision module and forward update mechanism. As illustrated in Figure 7, the integration of $BF$ with our proposed methods ($D+F$ and $S+F$) consistently improved accuracy across all model depths and both CIFAR10 and CIFAR100 datasets. Specifically, the $S+F$ configuration with $BF$ achieved peak performance of 93.27% on CIFAR10 and 72.02% on CIFAR100 with the VGG10 architecture, outperforming even the $BP$ baseline. In contrast, when $BF$ was applied to standard $PC$ or only with forward Update, we observed performance degradation rather than improvement in most cases. For models using only decaying precision or spiking precision without forward update, the effect of $BF$ was inconsistent and unpredictable across different network depths. These results suggest that the synergy between our proposed components is crucial—$BF$ appears to stabilize the training dynamics specifically when used in conjunction with both our energy balancing mechanisms (either decaying or spiking precision) and forward update. This interaction allows deeper networks to maintain stable gradients throughout the training process, resulting in more robust optimization and ultimately higher classification accuracy.

### v3

@@ D.1.2 Energy Propagation with Precision
Figure 5: Layer-wise Energy Distribution and Accuracy Comparison between PC and Decaying Precision/Spiking Precision with Forward Update in VGG5, VGG7 and VGG10 on the CIFAR10 dataset. The colored lines represent the total energy of the individual layers of the model. The vertical lines represent the train and test accuracies of the model.
Figure 6: Layer-wise Energy Distribution and Accuracy of Forward Update in VGG5, VGG7 and VGG10 on the CIFAR10 dataset. The colored lines represent the total energy of the individual layers of the model. The vertical lines represent the train and test accuracies of the model.
D.2 Forward Update (FU)
D.2.1 Neural activity divergency quantification
To quantify the neural activity divergence that Forward Update aims to solve, we conducted a new experiment on a VGG10 model trained on CIFAR-10. We measured the Mean Squared Error between the initial and final neural states, $MSE(x_{0}^{l},x_{T}^{l})$, at each weight update. We then calculated the ratio of the square root of this divergence to the energy used for the weight update in that layer. We term this metric the "Gap Ratio". As shown in Table 10, in the model trained without Forward Update, the Gap Ratio in the final layer (L10) is extremely large and unstable across epochs, indicating that the neural activity divergence completely dominates the weight update signal. This supports our hypothesis that this divergence causes errors to accumulate in the final layers, destabilizing learning. In contrast, the model trained with Forward Update (Table 11) shows a dramatically reduced and stable Gap Ratio in the final layer.
Table 10: Gap Ratio on VGG10/CIFAR10 trained without Forward Update.
w/o FU L1 L2 L3 L4 L5 L6 L7 L8 L9 L10 Train Acc. Test Acc.
Epoch 5 0.0 99.49 98.98 97.08 106.78 113.52 110.77 468.9 566.52 7065.72 73.75 74.83
Epoch 15 0.0 99.51 99.4 97.2 98.95 105.62 132.44 108.39 101.0 3807.9 51.86 51.9
Epoch 25 0.0 99.25 97.94 98.22 89.95 90.48 116.55 113.2 102.21 4716.88 61.21 62.3
Epoch 35 0.0 99.41 98.24 96.97 79.12 95.99 122.38 106.3 104.07 6498.41 60.31 60.14
Epoch 45 0.0 98.23 97.94 97.36 80.84 90.42 116.96 108.26 106.42 5362.02 59.87 61.42
Table 11: Gap Ratio on VGG10/CIFAR10 trained with Forward Update.
with FU L1 L2 L3 L4 L5 L6 L7 L8 L9 L10 Train Acc. Test Acc.
Epoch 5 0.0 292.58 112.77 224.2 120.89 346.42 143.95 248.32 100.78 5.14 88.27 85.66
Epoch 15 0.0 327.43 132.66 265.24 151.09 409.26 264.34 545.12 117.6 8.27 97.97 90.99
Epoch 25 0.0 323.0 135.82 256.2 149.77 407.2 269.14 619.95 115.07 9.43 99.55 92.08
Epoch 35 0.0 322.42 138.05 258.38 148.22 415.02 292.98 681.21 115.12 10.06 99.73 92.04
Epoch 45 0.0 323.25 135.72 261.67 149.45 417.04 312.4 731.19 115.87 10.54 99.83 92.22
To further isolate the effect of Forward Update, we performed an additional ablation where FU was applied only to the final three layers during the weight update, without any additional hyperparameter tuning. The results in Table 12 show that even this targeted application of Forward Update yields significant improvements in both training stability and final accuracy compared to the baseline. This reinforces that addressing the neural activity divergence in the deepest layers is a critical factor for success.
Table 12: Gap Ratio on VGG10/CIFAR10 when applying Forward Update only to the last 3 layers (L8, L9, L10).
L1 L2 L3 L4 L5 L6 L7 L8 L9 L10 Train Acc. Test Acc.
Epoch 5 0.0 96.21 57.22 50.78 32.03 53.59 30.16 49.77 66.91 42.19 72.75 74.13
Epoch 15 0.0 95.51 57.52 49.76 31.65 56.92 32.96 40.87 70.08 62.02 79.23 78.56
Epoch 25 0.0 95.58 59.17 46.63 16.04 91.18 34.19 39.63 73.17 67.89 81.4 80.14
Epoch 35 0.0 95.44 56.84 29.72 9.65 94.43 35.13 39.37 72.12 70.17 80.91 79.97
Epoch 45 0.0 95.51 56.88 37.85 15.11 93.49 34.8 38.59 71.82 72.16 80.02 79.4
D.2.2 Model Robustness Analysis
Predictive Coding (PC) networks are often noted for their inherent robustness compared to networks trained with Backpropagation Salvatori et al. (2021). Unlike our precision-weighting mechanisms, the Forward Update (F) method alters the core computational diagram of PC. Therefore, we conducted experiments to investigate whether this modification adversely affects model robustness.
Our first experiment provides a fair comparison on a VGG5 model, where standard PC and our PC+F variant achieve similar baseline accuracies on clean data. We trained both models on CIFAR-10 and evaluated their calibration under six types of data corruption across five levels of intensity. The range of the Adaptive Expected Calibration Error (adaECE) is reported in Table 13. Note that to ensure a meaningful ECE calculation, we scale the output logits before applying the softmax function. The results indicate that our PC+F model maintains a robustness profile that is comparable to, and at higher corruption levels, superior to that of standard PC.
Table 13: AdaECE range at different levels of corruption using VGG5.
Corruption Level PC+FU PC
0.1 [0.091665, 0.327372] [0.037817, 0.388640]
0.2 [0.089805, 0.349025] [0.036528, 0.406428]
0.3 [0.034430, 0.363157] [0.035065, 0.419023]
0.4 [0.021906, 0.375255] [0.029735, 0.428372]
0.5 [0.026589, 0.383087] [0.033315, 0.434768]
Furthermore, to evaluate our methods on deeper models, we compared ourPC+S+FU model against the BP baseline on the VGG10 architecture. As shown in Table 14, our method exhibits a robustness profile that is highly comparable to backpropagation across all tested corruption intensities. Taken together, these experiments demonstrate that our proposed methods not only enable the training of deeper PCNs but do so while preserving the desirable property of model robustness.
Table 14: AdaECE range at different levels of corruption using VGG10.
Corruption Level BP PC+S+FU
0.1 [0.039071, 0.485300] [0.039269, 0.486342]
0.2 [0.043376, 0.490541] [0.045973, 0.486978]
0.3 [0.043943, 0.494404] [0.047871, 0.486432]
0.4 [0.067549, 0.496167] [0.077340, 0.486966]
0.5 [0.132682, 0.499181] [0.143458, 0.486497]
D.2.3 On the Biological Plausibility of Forward Update
A core motivation for using PCNs over backpropagation is their biological plausibility — particularly their use of local learning rules and temporally local computations. The Forward Update (F) mechanism, while effective, seems to introduce non-locality in time by requiring each synapse to store its initial feedforward activity $\mu_{0}^{l}$ until convergence.
However, the description of FU in our manuscript was chosen for conceptual clarity; it is not a fundamental requirement of the method. In practice, the initial feed-forward state can be re-computed through a temporally local and biologically plausible process. This is achieved by introducing a "free relaxation" phase after the inference learning and before the weight update. In this phase, the label clamp is removed, and the network settles to a new equilibrium with only the sensory input clamped, just as in Equilibrium Propagation (Ernoult et al., 2020) before the nudging phase.
In this setting, the network naturally converges to the state corresponding to its feed-forward prediction (Frieder & Lukasiewicz, 2022). Since the weights have not yet been updated, this re-computed state is identical to the $\mu_{0}^{l}$ in our formulation. This eliminates the need for long-term storage and resolves the concern of temporal non-locality. The primary contribution of our F method is that it identifies and solves a critical failure mode in deep predictive coding networks: the accumulation of errors caused by the divergence of neural activities from their initial predictions. While the biologically plausible implementation we describe requires extra computation, future work can focus on developing mechanisms that are both fully plausible and computationally efficient for PC.
D.3 Ablation study on Forward Update and Spiking precision
Table 15: Test accuracies of different algorithms without BatchNorm Freeze across datasets and architectures.
Dataset Algorithm VGG5 VGG7 VGG10 ResNet10 ResNet18
CIFAR10 BP $89.43^{\pm 0.12}$ $89.91^{\pm 0.12}$ $92.21^{\pm 0.08}$ $92.21^{\pm 0.20}$ $\mathbf{92.32^{\pm 0.22}}$
PC $87.98^{\pm 0.11}$ $84.62^{\pm 0.10}$ $72.75^{\pm 6.03}$ $69.85^{\pm 2.12}$ $15.63^{\pm 7.22}$
iPC $85.51^{\pm 0.12}$ $80.15^{\pm 0.18}$ $63.83^{\pm 0.33}$ $62.34^{\pm 0.27}$ $21.90^{\pm 1.51}$
PC+Spiking Precision (S) $88.06^{\pm 0.16}$ $82.16^{\pm 0.14}$ $78.09^{\pm 0.61}$ $80.61^{\pm 0.20}$ $80.07^{\pm 0.21}$
PC+Forward Update (F) $88.79^{\pm 0.04}$ $87.43^{\pm 0.30}$ $87.33^{\pm 0.14}$ $66.94^{\pm 0.76}$ $25.50^{\pm 3.12}$
iPC+S $89.32^{\pm 0.13}$ $\mathbf{90.25^{\pm 0.06}}$ $\mathbf{93.03^{\pm 0.18}}$ $\mathbf{92.39^{\pm 0.04}}$ $91.96^{\pm 0.07}$
PC+S+F $\mathbf{89.45^{\pm 0.18}}$ $90.08^{\pm 0.21}$ $92.07^{\pm 0.10}$ $92.04^{\pm 0.04}$ $91.91^{\pm 0.09}$
CIFAR100 (Top-1) BP $66.28^{\pm 0.23}$ $65.36^{\pm 0.15}$ $69.35^{\pm 0.16}$ $69.23^{\pm 0.09}$ $\mathbf{71.46^{\pm 0.12}}$
PC $60.00^{\pm 0.19}$ $56.80^{\pm 0.14}$ $45.86^{\pm 1.70}$ $27.62^{\pm 3.03}$ $1.59^{\pm 0.02}$
iPC $56.07^{\pm 0.16}$ $43.99^{\pm 0.30}$ $1.37^{\pm 0.37} 22$$2.91^{\pm 0.23}$ $1.53^{\pm 0.06}$
PC+Spiking Precision (S) $59.18^{\pm 0.20}$ $56.98^{\pm 0.19}$ $51.56^{\pm 0.16}$ $50.23^{\pm 0.20}$ $22.92^{\pm 0.15}$
PC+Forward Update (F) $65.34^{\pm 0.07}$ $64.50^{\pm 0.14}$ $61.69^{\pm 0.79}$ $39.89^{\pm 0.90}$ $3.42^{\pm 0.10}$
iPC+S $65.54^{\pm 0.62}$ $65.76^{\pm 0.12}$ $\mathbf{69.84^{\pm 0.17}}$ $\mathbf{70.02^{\pm 0.24}}$ $70.38^{\pm 0.20}$
PC+S+F $\mathbf{66.49^{\pm 0.15}}$ $\mathbf{66.34^{\pm 0.22}}$ $69.08^{\pm 0.08}$ $68.99^{\pm 0.18}$ $70.81^{\pm 0.08}$
[…]
PC $84.97^{\pm 0.19}$ $83.00^{\pm 0.09}$ $74.61^{\pm 1.08}$ $57.93^{\pm 2.62}$ $5.89^{\pm 0.12}$
iPC $78.91^{\pm 0.23}$ $73.23^{\pm 0.30}$ $48.35^{\pm 0.79}$ $46.41^{\pm 0.31}$ $6.33^{\pm 0.26}$
PC+Spiking Precision (S) $84.58^{\pm 0.12}$ $83.61^{\pm 0.15}$ $8.62^{\pm 0.15} 77$$7.84^{\pm 0.23}$ $53.89^{\pm 0.06}$
PC+Forward Update (F) $85.48^{\pm 0.08}$ $84.05^{\pm 0.07}$ $76.73^{\pm 0.92}$ $69.48^{\pm 0.70}$ $15.25^{\pm 0.04}$
[…]
PC+S+F $\mathbf{86.36^{\pm 0.11}}$ $4.53^{\pm 0.15} 88$$66^{\pm 0.14} 88.$$8.45^{\pm 0.16}$ $\mathbf{90.47^{\pm 0.11}}$
To dissect the individual and combined contributions of our primary algorithmic modifications, Spiking Precision (S) and Forward Update (F), we conducted a detailed ablation study, the results of which are presented in Table 15. This analysis systematically evaluates each component’s impact on both standard PC and iPC across various architectures and datasets, excluding the effects of BatchNorm Freezing to isolate the core mechanisms.
The results reveal a clear and complementary relationship between Spiking Precision and Forward Update for standard PC. When applied in isolation, Forward Update (PC+F) significantly improves performance on VGG-style architectures, stabilizing training and preventing the sharp accuracy degradation seen in the baseline PC as depth increases. For instance, on CIFAR10, PC+F maintains an accuracy of around $87\%$ on VGG10, whereas the baseline PC drops to $72.75\%$. However, Forward Update alone is insufficient for training deep residual networks; its performance on ResNet18 is only marginally better than the baseline, failing to overcome the catastrophic failure. This suggests that while F effectively mitigates weight update divergence, it does not solve the underlying problem of energy imbalance in architectures with skip connections.
Conversely, Spiking Precision alone (PC+S) offers a substantial improvement on ResNet models, preventing the complete collapse of training. On ResNet18 with CIFAR10, it achieves an accuracy of $80.07\%$, a dramatic recovery from the baseline’s $15.63\%$. This confirms its crucial role in rebalancing energy and ensuring a viable error signal reaches the early layers in models with skip-connections. However, on its own, it does not elevate performance to the level of backpropagation.
The true strength of our approach is demonstrated when both components are combined. The PC+S+F model consistently achieves performance on par with, and occasionally exceeding, backpropagation across all tested architectures, including the challenging ResNet18. This powerful synergy underscores that both mechanisms are essential: Spiking Precision addresses the signal propagation problem, while Forward Update addresses the update stability problem.
Interestingly, for iPC, the addition of Spiking Precision alone (iPC+S) is sufficient to achieve state-of-the-art performance, rivaling both BP and the fully-equipped PC+S+F. This indicates that the incremental nature of iPC, where weights are updated at every inference timestep, inherently prevents the large divergence between forward neural state and backward neural states that Forward Update is designed to correct. With its continuous adaptation, iPC only requires the energy rebalancing provided by Spiking Precision to successfully train deep architectures.
D.4 BatchNorm Freezing (BF)
To isolate the precise source of instability when applying BatchNorm (BN) to Predictive Coding Networks (PCNs), we sought to determine whether the problem stems from the learnable affine parameters ($\gamma,\beta$) or from the iterative updating of batch statistics ($\mu_{B},\sigma^{2}_{B}$) during the inference phase. While the affine parameters can be a source of overfitting in standard training, we hypothesized that the unique, multi-step inference process of PCNs creates a different challenge: the repeated processing of a single mini-batch causes the batch statistics themselves to overfit, destabilizing the network dynamics. Our BatchNorm Freezing (BF) method is designed specifically to solve this issue.
To test this hypothesis, we conducted an ablation study on the VGG10-CIFAR10 task. The results presented in Table 16, provide clear evidence for our claim. Removing the affine parameters from standard BN still had a negligible effect on the performance degradation ($24.31\%$ and $24.28\%$), indicating that these parameters are not the source of the problem. In contrast, freezing the statistics during inference improved performance over the without batch normalization baseline. This confirms that the iterative updates to batch statistics are the primary cause of instability and that our proposed BF method is an effective solution.
Table 16: Ablation study on different BatchNorm strategies (VGG10-CIFAR10).
Method Test Accuracy (%)
Without BN $92.07\pm 0.10$
Standard BN $24.31\pm 4.51$
BN without affine $24.28\pm 3.19$
BF without affine $93.18\pm 0.06$
BF $\mathbf{93.27\pm 0.10}$
D.4.1 Ablation study on batchnorm Freezing
@@
Our investigation reveals that BatchNorm Freezing ($BF$) significantly enhances model performance when combined with our precision module and forward update mechanism. As illustrated in Figure 7, the integration of $BF$ with our proposed methods ($PC+S+F$) consistently improved accuracy across all model depths and both CIFAR10 and CIFAR100 datasets. Specifically, the $PC+S+F$ configuration with $BF$ achieved peak performance of 93.27% on CIFAR10, 72.02% on CIFAR100 with the VGG10 architecture and 71.3% on CIFAR100 with the ResNet10 architecture, outperforming even the $BP$ baseline. In contrast, when $BF$ was applied to standard $PC$ or only with forward Update, we observed performance degradation rather than improvement in most cases, the effect of $BF$ was inconsistent and unpredictable across different network depths. These results suggest that the synergy between our proposed components is crucial—$BF$ appears to stabilize the training dynamics specifically when used in conjunction with both our energy balancing mechanisms and forward update. This interaction allows deeper networks to maintain stable gradients throughout the training process, resulting in more robust optimization and ultimately higher classification accuracy.
The effect of BF on iPC-based models is more nuanced. While it improves performance on shallower architectures, a slight performance degradation is observed in deeper models. This discrepancy can be attributed to the nature of iPC’s update rule. Because iPC updates weights and neural activities simultaneously, updating BN’s running statistics requires an additional, separate weight update step. This modification alters the computational graph compared to the baseline model (without BF) that was used for the hyperparameter search, which could explain the suboptimal performance in more complex architectures.
Appendix E Hyperparameter Analysis
E.1 Hyperparameter Importance
To assess the sensitivity of our approach to different hyperparameters, we conducted a hyperparameter importance analysis using a functional ANOVA (fANOVA) based method. This score quantifies the contribution of each hyperparameter to the final optimal validation accuracy. The results for CIFAR10 and CIFAR100 on various architectures are summarized in Tables 17 and 18. The analysis shows that while the learning rate ($lr_{w}$) is consistently the most critical parameter, our new parameters (like $\alpha/k$ in Precision) are not overly sensitive and show stable influence across settings, indicating a desirable robustness in our proposed methods.
Table 17: Hyperparameter Importance Scores (%) on CIFAR-10.
Arch. Method Activation T $\alpha/k$ $lr_{x}$ $lr_{w}$ $momentum_{x}$ $weight\_decay_{w}$
VGG5 BP 0.94 98.99 0.07
PC 5.96 1.00 77.29 10.98 3.51 1.26
iPC 2.20 1.15 1.53 88.75 1.83 3.25
iPC+S 0.85 0.52 0.22 1.13 94.02 0.31 2.94
PC+S+F 1.68 1.63 0.53 6.64 80.44 8.19 0.89
PC+D+F 6.34 0.50 0.12 11.53 74.62 2.80 4.08
VGG7 BP 25.94 61.56 12.49
PC 15.65 0.57 54.54 23.56 4.35 1.33
iPC 1.68 1.63 6.64 80.44 8.19 0.89
iPC+S 1.68 1.63 0.89 6.64 80.44 8.19 0.89
PC+S+F 3.30 0.28 4.13 10.29 54.20 1.80 25.99
PC+D+F 14.78 3.05 0.01 9.00 67.45 1.17 4.54
VGG10 BP 9.69 87.86 2.45
PC 48.10 5.62 4.02 16.15 25.32 0.79
iPC 0.08 0.48 34.55 58.50 4.73 1.66
iPC+S 4.85 1.80 1.72 15.15 67.61 5.64 3.22
PC+S+F 4.84 1.55 2.08 9.80 77.14 2.16 2.43
PC+D+F 8.25 2.79 0.04 1.74 75.95 9.01 2.22
ResNet10 BP 6.60 93.30 0.10
PC 8.87 0.46 30.41 33.60 21.77 4.89
iPC 1.47 1.31 19.18 61.87 2.87 13.30
iPC+S 4.74 0.22 1.07 11.68 59.33 11.79 11.16
PC+S+F 7.86 2.04 17.93 9.63 57.00 4.32 1.22
ResNet18 BP 7.80 92.00 0.20
PC 11.39 2.86 55.11 12.75 8.48 9.40
iPC 3.76 1.57 14.85 47.67 11.08 21.07
iPC+S 5.82 0.55 1.95 2.95 78.29 8.83 1.61
PC+S+F 1.73 17.97 5.98 16.95 48.95 3.05 5.36
Table 18: Hyperparameter Importance Scores (%) on CIFAR-100.
Arch. Method Activation T $\alpha/k$ $lr_{x}$ $lr_{w}$ $momentum_{x}$ $weight\_decay_{w}$
VGG5 BP 9.49 90.50 0.01
PC 3.10 3.03 83.65 6.56 2.55 1.12
iPC 1.14 2.18 1.37 91.29 3.04 0.99
iPC+S 3.37 0.62 5.74 2.92 86.36 0.79 0.20
PC+S+F 7.12 8.64 1.09 1.85 75.10 5.47 0.73
PC+D+F 10.43 4.03 0.07 29.70 52.92 1.50 1.35
VGG7 BP 3.73 92.30 3.97
PC 9.12 1.62 68.17 10.75 7.14 3.20
iPC 3.26 0.64 1.87 57.30 36.54 0.40
iPC+S 13.47 0.37 5.47 9.96 61.09 4.30 5.34
PC+S+F 2.38 5.03 0.52 7.41 79.25 3.43 1.97
PC+D+F 6.98 0.49 0.03 0.43 86.89 0.02 5.17
VGG10 BP 1.88 98.10 0.02
PC 22.79 3.60 6.17 62.83 3.88 0.74
iPC 6.01 2.51 4.48 40.86 43.68 2.47
iPC+S 4.12 0.04 6.89 5.75 77.72 3.99 1.50
PC+S+F 3.76 0.40 1.25 6.71 82.10 3.04 2.74
PC+D+F 1.34 0.39 1.70 4.41 79.06 3.24 9.87
ResNet10 BP 8.67 90.23 1.10
PC 20.05 4.46 10.44 46.35 14.47 4.23
iPC 11.93 3.77 9.98 59.41 13.44 1.47
iPC+S 0.32 2.10 6.33 21.00 44.40 1.57 24.28
S+F 1.93 0.53 1.23 23.06 66.37 5.77 1.12
ResNet18 BP 1.49 98.21 0.30
PC 0.84 4.17 24.00 42.24 7.79 20.96
iPC 2.37 7.89 13.94 19.73 9.87 46.20
iPC+S 1.20 0.42 0.78 52.00 41.70 1.90 1.99
S+F 1.30 0.46 0.79 14.49 34.98 45.45 2.54
E.2 Hyperparameter Transferability
To evaluate if hyperparameter searching is required for each setting, we conducted two sets of experiments to evaluate hyperparameter transferability.
Across Datasets: We took the optimal hyperparameters found on CIFAR-10 and applied them to CIFAR-100, and vice versa, for the VGG7 architecture.
Across Architectures: We took the optimal hyperparameters from VGG5 and VGG7 and applied them to the VGG10 model on CIFAR-10. Since the hyperparameter T needs to be larger than the number of layers, the T we used in these experiments is $max(10,T_{optimal})$.
The results, presented in Table 19 and 20, suggest that while optimal performance requires dedicated tuning, the hyperparameters show a reasonable degree of transferability, especially for our proposed methods. This indicates they are not pathologically sensitive to the specific dataset or architecture.
Table 19: Test Accuracies from Hyperparameters (HPs) Transfer Across Datasets on VGG7.
Method On CIFAR-10 On CIFAR-100
Optimal CIFAR-10 HPs Optimal CIFAR-100 HPs Optimal CIFAR-100 HPs Optimal CIFAR-10 HPs
BP $9.91^{\pm 0.12} 88$$8.75^{\pm 0.07}$ $65.36^{\pm 0.15}$ $65.23^{\pm 0.24}$
PC $84.62^{\pm 0.10}$ $78.38^{\pm 0.17}$ $56.80^{\pm 0.14}$ $51.65^{\pm 0.28}$
S+F $90.08^{\pm 0.21}$ $89.54^{\pm 0.09}$ $66.34^{\pm 0.22}$ $63.66^{\pm 0.06}$
iPC $80.15^{\pm 0.18}$ $74.35^{\pm 0.90}$ $43.99^{\pm 0.30}$ $37.26^{\pm 0.07}$
iPC+S $90.25^{\pm 0.06}$ $90.14^{\pm 0.20}$ $65.76^{\pm 0.12}$ $61.56^{\pm 1.56}$
Table 20: Test Accuracies from Hyperparameters (HPs) Transfer Across Architectures (on CIFAR-10).
Method VGG10 with Optimal HPs VGG10 with VGG7 HPs VGG10 with VGG5 HPs
BP $92.21^{\pm 0.08}$ $90.70^{\pm 0.14}$ $90.67^{\pm 0.22}$
PC $72.75^{\pm 6.03}$ $79.20^{\pm 0.19}$ $49.66^{\pm 0.54}$
S+F $92.07^{\pm 0.10}$ $88.95^{\pm 0.18}$ $90.67^{\pm 0.11}$
iPC $63.83^{\pm 0.33}$ $72.30^{\pm 0.55}$ $73.78^{\pm 0.15}$
iPC+S $93.03^{\pm 0.18}$ $90.60^{\pm 0.60}$ $92.29^{\pm 0.13}$